A glass box 7 x 12 x 18 cm, closed on all six sides, is partially filled with colored water. When the box is placed on one of its 7 x 12 sides, the water level is 15 cm above the table. If the box is placed on one of its 7 x 18 sides, what will be the water level above the table,in centimeters?
Answer by Theo(13342)

the glass box is 7*12*18 cm

when placed on 7 * 12 side, volume of the water is 7 * 12 * 15 = 1260 cubic cm.

when placed on 7 * 18 side, the volume of the water is still 1260 cubic cm.

let x = the height.

you get:

l * w * h = volume which becomes:

7 * 18 * x = 1260

solve for x to get x = 1260/(7*18) = 10 inches high.

that's your answer.